OpenStudy (anonymous):

Can someone help me with this algebra 2 assignment? I'm lost & I would appreciate it so much!?

OpenStudy (anonymous):

1. Write a word problem describing the time it takes to complete an activity individually and with a friend. For example, if John takes 2 hours to mow his lawn and it takes his sister Maria 4 hours to mow the same lawn, how long would it take John and Maria to mow the lawn together?

OpenStudy (anonymous):

2. Write a rational equation based upon the word problem you created. The rational equation based upon the scenario above would be 1 over 2 + 1 over 4 = 1 over x. Each fraction represents the amount of the lawn mowed in one hour. The fraction of 1 over 2 is John’s portion. Maria’s share is represented by 1 over 4. The time it would take for both of them to mow the lawn is represented by 1 over x. 3. Solve the rational equation. Show your work.

OpenStudy (firejay5):

Here is a similar example for #1: https://answers.yahoo.com/question/index?qid=20101213170345AAT0toR three workers can do a contraction in 5 hours a hero worker joins their group and they finish it in 2.5 hours how much would it take for the super worker to finish the contraction alone? x= amount of work one worker does in one hour y= amount of work super hero worker does in an hour 3*x*5=15x=the amount of whole contraction 3*x*2.5+y*2.5=also the amount of contraction so we can write : 15x=7.5x+2.5y => y=3x and if t is the time it takes for the hero to finish the job

OpenStudy (anonymous):

Okay so does that mean for the first part I have to make up a word problem?

OpenStudy (anonymous):

@Firejay5 Okay so does that mean for the first part I have to make up a word problem?

OpenStudy (firejay5):

do exactly what the problem is presented with and #1 should be easy.

OpenStudy (anonymous):

@Firejay5 I'm going to make up a word problem so... Two guys can finish a house painting in 4 hours One guy comes along & they finish it in 2.5 hours How long would it take for the one guy to finish the house painting by himself?

OpenStudy (anonymous):

@Firejay5 I get the making up the word problem part but I need help with the actual work

OpenStudy (anonymous):

@Firejay5 since i said 4 hours it would probably go something like this... the letter x would equal how long it takes the two guys to paint the house in one hour the letter y would equal how long it takes the one to paint the house in an hour right?